Write each equation in slope-intercept form of the equation of a line. Underline the slope and circle the y-intercept in each equation.
3x+4y–12=0
y - 3= - 3x/4
The slope is the coefficient of x, that is slope = -3/4. The y intercept is the value of y when x - 0. Plugging back x = 0 in the equation above, you see that y = 3.
It's always a good idea to check directly in the original equation. 3*0 + 4*3 - 12 = 0

Answer by t0hierry(194)

3x+4y–12=0
Solve for y
y = - 3x/4 + 3
The slope is the coefficient of x, that is slope = -3/4. The y intercept is the value of y when x - 0. Plugging back x = 0 in the equation above, you see that y = 3.
It's always a good idea to check directly in the original equation. 3*0 + 4*3 - 12 = 0 as expected